# Quantum noise spectroscopy of superconducting critical dynamics and vortex fluctuations in a high-temperature cuprate

### v1

@@ Abstract
Characterizing the low-energy dynamics of quantum materials is crucial to our understanding of strongly correlated electronic states. However, extracting universal dynamical features requires resolving correlations at both low energy and momentum. Here, we introduce nitrogen-vacancy (NV) centers in diamond as a novel and powerful quantum sensing platform of superconducting materials. We demonstrate the strengths of our approach by probing several low-energy phenomena in high-$T_{c}$ cuprate Bi2Sr2CaCu2O8+δ (BSCCO) — gapless quasiparticle excitations, critical fluctuations at the metal-superconductor transition and kinetics of vortices. In the absence of an applied magnetic field, we find a sharp reduction in the NV relaxation time ($T_{1}$) near the critical temperature $T_{c}\approx 90~{}$K, attributed to supercurrent-fluctuation induced magnetic noise. Crucially, the temperature-scaling of the noise near criticality deviates from the Bardeen–Cooper–Schrieffer (BCS) mean-field prediction and reflects critical order parameter fluctuations, allowing us to determine both static and dynamical critical exponents for the transition. When a small field is applied, we detect a broad and asymmetric reduction of $T_{1}$ near $T_{c}$, indicating significant field-induced smearing of the transition. By analyzing the scaling of the BSCCO-induced relaxation rate with the field strength, we unveil evidence in favor of a vortex liquid phase. Finally, deep inside the superconducting phase, we employ NV decoherence ($T_{2}$) spectroscopy to observe strong magnetic fluctuations in the low-frequency regime, suggesting the presence of complex vortex-solid fluctuations. Our results establish NV-based noise spectroscopy as a versatile platform for probing dynamical phenomena in superconductors, with frequency and length scales complementary to existing techniques.
@@
Introduction — Strong electronic interactions are responsible for a plethora of complex emergent phenomena observed in quantum materials. This is perhaps best exemplified by unconventional superconductors, where strong electronic interactions induce a macroscopic quantum coherence of electron-pairs and dissipationless flow of currents tinkham ; keimer2015quantum ; zhou2021high . Key features of this state are encoded in its low-energy dynamics, manifesting as quasiparticle excitations arising from the breaking of electron pairs bardeen1957microscopic ; bardeen1957theory ; bogoljubov1958new , slow collective modes such as critical pairing fluctuations near the metal–superconductor phase transition larkin2005theory ; schmid1966time ; maki1968critical ; cyrot1973ginzburg ; schuller2006time , or topological excitations such as vortices carrying quantized magnetic flux tinkham ; blatter1994vortices ; halperin1979resistive ; Beasley ; Nelson_VL . Over the years, several experimental techniques have been developed to probe these universal, low-energy features comin2016resonant ; kaindl2000ultrafast ; de2021colloquium ; parker:2010 ; essaleh:2018 ; hore:2015 .
@@
Figure 1: NV noise spectroscopy of superconducting dynamics. (a) Summary of different experimental techniques for probing superconductivity, organized by their frequency and lengthscales, highlighting the complementarity of our qubit-based noise spectroscopy (NV $T_{1}$ and $T_{2}$) and DC magnetometry barry:2020b . Techniques depicted include: electrical transport essaleh:2018 , superconducting quantum interference device (SQUID) kirtley:1999 ; dusad:2019 , Lorentz microscopy harada:1992 ; silhanek:2021 , X-ray/neutron scattering comin2016resonant ; muhlbauer:2019 , magnetic resonance with nuclear, electronic, or muonic spins (NMR/ESR/µSR) hore:2015 ; hillier:2022 , angle-resolved photoemission spectroscopy (APRES) liu2008development ; kitamura2022development , low-energy and transmission electron microscopy (LEEM/TEM) altman:2010 ; moradifar:2023 , and scanning tunneling microscope (STM) parker:2010 . Some techniques, while capable of accessing high momentum data, require macroscopic sample volumes and, thus, are only suitable for probing mm-scale samples. Techniques like pump-probe spectroscopy kaindl2000ultrafast ; de2021colloquium , which are known to drive the sample strongly out of equilibrium, are not included. (b) Schematic of the experimental set-up: a thin film ($\sim 200$ nm thickness) of BSCCO is directly exfoliated onto a diamond plate containing an ensemble of NV centers within a thin layer ($\sim 50$ nm) beneath its surface. Current fluctuations in the BSCCO sample induce magnetic field noise at the NV locations which can be detected by either depolarization ($T_{1}$) or decoherence ($T_{2}$) noise spectroscopy. An out of plane magnetic field $H$ is applied, which is parallel to the quantization axis of the [111] NV group. The NV exhibits a splitting of its $|m_{s}=\pm 1\rangle$ sublevels, which is determined by the local magnetic field $B_{z}$, centered around a zero-field splitting $D_{\mathrm{gs}}=2.87~{}$GHz. (c) Schematic H-T phase diagram of a typical type-II superconductor tinkham ; keimer2015quantum ; zhou2021high . The green arrows indicate zero-field cooling ($H=0$) and field-cooling ($H>0$) pathways used in our experiments.
@@
Broadly, techniques for measuring correlated quantum dynamics may be categorized into scattering methods, local spectroscopic probes, and bulk measurements [Fig. 1(a)]. Scattering probes (e.g. X-ray scattering comin2016resonant ) provide both energy and momentum-resolved data, but they typically require experimental apparatus beyond table-top settings, large sample volumes, and can drive the system out of equilibrium kaindl2000ultrafast ; de2021colloquium . Local probes such as scanning tunneling microscopy (STM) parker:2010 are applicable to low-dimensional samples and provide lattice-scale spatial resolution, yet they exhibit limited momentum resolution and typically operate at low temperatures. By contrast, bulk measurements (including transport essaleh:2018 and nuclear magnetic resonance hore:2015 ) can operate over a wide temperature range, but generally demand large sample volume. Therefore, resolving spectral features at long wave-lengths below the meV scale in two-dimensional materials and thin film samples remains a challenge, leading to an unfulfilled gap for experimental techniques that can probe equilibrium dynamics at both low-energies and a variety of length-scales.
@@
In this work, we fill this gap by introducing a non-invasive, table-top experimental technique — quantum noise spectroscopy — to probe superconducting dynamics with simultaneous spatial and temporal resolution. Specifically, by using an ensemble of nitrogen-vacancy (NV) centers positioned in close proximity to a thin film high-$T_{c}$ cuprate — Bi2Sr2CaCu2O8+δ (BSCCO, Bi-2212), we perform noise spectroscopy of superconducting fluctuations as a function of temperature, frequency and applied magnetic field. We demonstrate in-situ detection of spatially resolved dynamics spanning a broad frequency range (MHz to GHz) and length-scales (nm to $\mathrm{\mu m}$), thus offering a powerful complementary toolset to existing techniques for probing superconductivity [Fig. 1(a)].
@@
Our main results are threefold. First, at zero applied field, we study the fluctuations arising from both long-lived quasiparticle excitations at low temperature, and collective critical modes near the metal-superconductor transition temperature $T_{c}$. These fluctuations induce GHz-scale magnetic noise at the location of the nearby NVs which can be characterized via NV depolarization ($T_{1}$) spectroscopy. Leveraging this capability, we quantitatively measure the characteristic time-scale $\tau_{0}$ of critical slowing down of pairing fluctuations near $T_{c}$, as well as the static and dynamical critical exponents associated with the phase transition landau2013statistical ; landau1980statistical ; hohenberg1977theory .
@@
Second, we directly probe the broadening of the transition in the presence of a small applied field $H$. At non-zero $H$, the system hosts superconducting vortices whose dynamics induce additional magnetic field fluctuations. The field dependence of the NV relaxation rate is consistent with the existence of an intermediate-temperature vortex liquid phase blatter1994vortices ; halperin1979resistive , where magnetic flux lines diffuse through the superconducting film.
@@
Finally, we employ decoherence ($T_{2}$) noise spectroscopy to study lower-frequency fluctuations in the MHz regime machado2023quantum . Deep within the superconducting phase, we observe a distinct noise mechanism that significantly enhances the NV decoherence rate. These strong, low-frequency magnetic fluctuations potentially arise from thermally assisted flux motion in a vortex solid phase blatter1994vortices ; MFisher89 ; fisherfisherhuse .
@@
Figure 2: Observing Meissner effect and superconducting vortices. (a) ODMR spectrum of NV centers under BSCCO with an applied field $H=15$ G at temperatures from 10 K to 120 K. The colors represent the normalized fluorescence contrast (FL) of the spectrum. The gray dashed line marks the superconducting transition temperature $T_{c}\approx 90$ K. Top: at 10 K, ODMR spectrum of NV centers under BSCCO at $H=0$ (yellow), where the intrinsic splitting $2\Pi_{\perp}$ originates from the local electric field mittiga2018imaging . At $H=15$ G, the spectrum (blue) remains largely unchanged, indicating a strong magnetic field repulsion due to the Meissner effect. Bottom: when the bias field is not suppressed (120 K), the [111] NV group aligned with the bias field exhibits a splitting $\Delta$, while the remaining NVs in other three crystal orientations are degenerate, contributing to the inner two peaks. (b) Optical image of the thin film BSCCO exfoliated on a diamond surface. (c) Local magnetic field $B_{z}$ as a function of the applied field $H$ along a line-cut across the BSCCO [black dashed line in (b)] at 30 K. Compared to NV centers far away from BSCCO, those under BSCCO (gray region) experience smaller local magnetic fields due to the Meissner suppression. Near the edge of the BSCCO, $B_{z}$ gets enhanced due to the concentration of the magnetic field lines. (d) Schematic of magnetic flux penetrating a superconductor in its vortex phase. Top: local field image of a defect-trapped superconducting vortex at $H=40$ G and $T=40$ K. (e) Schematic of magnetic flux being repelled by a superconductor in its Meissner phase.
@@
Experimental setup— For our experiments, we choose Bi2Sr2CaCu2O8+δ (Bi-2212), a prominent type-II superconductor with a layered crystal structure and a high critical temperature $T_{c}$ exceeding the boiling point of liquid nitrogen keimer2015quantum ; zhou2021high ; blatter1994vortices . This compound exhibits a rich and complex H-T phase diagram, making it an ideal platform for investigating a variety of different types of low-energy dynamics [Fig. 1(c)]. Bulk crystals of nearly optimally doped BSCCO-2212 are grown by the traveling floating zone method wen2008large , with $T_{c}\approx 90~{}$K measured via transport (see Methods). To perform noise spectroscopy with NV centers, we exfoliate the BSCCO to a 200200200~{}200nm thick film and transfer it directly onto a diamond sample containing an ensemble of NV centers in a thin layer $\sim 50$ nm beneath the diamond surface [Fig. 1(b)].
Figure 3: Zero-field superconducting fluctuations. (a) $T_{1}$ decay profile for NV centers located under BSCCO (left, blue) and away from BSCCO (right, orange) at 110, 93 and 40 K. Inset: Location of the measured NV centers with respect to the BSCCO sample. (b) The $T_{1}$ relaxation timescales as a function of temperature measured on the NV centers located under and away from BSCCO. (c) Spin relaxation rate, $1/T_{1}$, along a spatial line-cut across the BSCCO sample [see (a) inset] at 110, 93 and 40 K. The relaxation rate increases for NV centers under BSCCO (shadow region), particularly near the critical temperature $T_{c}$. (d) The BSCCO-induced decay rate of NVs, $\Gamma_{1}(T)$, as a function of temperature with $H=0$. Inset: a power-law relationship at low temperatures: $\Gamma_{1}(T)\propto T^{2}$, is consistent with a mean-field approach accounting for nodal quasiparticles in d-wave superconductors. (e) $\Gamma_{1}(T)$ as a function of the absolute temperature deviation from the critical temperature, $|T-T_{c}|$, for both the metallic and superconducting (SC) sides of the transition. BCS mean-field quasiparticle dynamics (dashed line) predicts $\Gamma_{1}(T)\propto|T-T_{c}|^{-x}$ with $x=1/2$. However, our experimental data is consistent with $x\approx 1$, suggesting that additional superconducting order parameter fluctuations are important. This data aligns closely with a theoretical model of critical Langevin dynamics of the order parameter (solid lines).
@@
Each NV center hosts a spin-1 electronic ground state, with Hamiltonian rovny2024nanoscale :
@@
where {$S_{x}$, $S_{y}$, $S_{z}$} are the spin-1 operators with the N–V axis defining the spin-quantization axis ($\hat{z}$), $D_{\mathrm{gs}}=2.87~{}$GHz is the zero-field splitting between $|m_{s}=0\rangle$ and $|m_{s}=\pm 1\rangle$ sublevels, $\Pi_{\perp}=\sqrt{\Pi_{x}^{2}+\Pi_{y}^{2}}=6.6~{}$MHz characterizes the coupling to local electric fields mittiga2018imaging , $B_{z}$ is the local magnetic field along the NV axis, and $\gamma_{e}$ is the electronic spin gyromagnetic ratio.
@@
The NV energy levels can be probed via optically detected magnetic resonance (ODMR) spectroscopy where one detects a decrease in NV fluorescence intensity when sweeping the frequency of an applied microwave through the resonances $|m_{s}=0\rangle\leftrightarrow|m_{s}=\pm 1\rangle$. In the experiment, we use a [111]-cut diamond crystal and apply a tunable magnetic field $H$ applied along the out-of-plane direction [Fig. 1(b)]. In this geometry, the NV crystallographic group aligned with the applied field exhibits the largest splitting, $\Delta=\sqrt{(2\Pi_{\perp})^{2}+(2\gamma_{e}B_{z})^{2}}$ between $|m_{s}=\pm 1\rangle$ sublevels, where $B_{z}$ includes both contributions from the applied field $H$ and the $B^{\mathrm{s}}_{z}$ generated from the superconducting sample ($B_{z}=H+B_{z}^{\mathrm{s}}$). The other three NV groups are always observed to be degenerate.
@@
Imaging the Meissner effect and flux vortices — Using the NV centers as sensitive local magnetometers, we begin by determining the presence of superconductivity in BSCCO via the DC Meissner effect schlussel2018wide ; yip2019measuring ; lesik2019magnetic ; bhattacharyya2023imaging ; nusran2018spatially ; ho2024studying . After zero-field cooling the sample down to 101010~{}10K (well below $T_{c}$), we perform an ODMR measurement with $H=0$ [Fig. 2(a)]. The resulting spectrum shows two resonances separated by the intrinsic splitting $2\Pi_{\perp}$. When we then apply a small field $H=15~{}$G, the ODMR spectrum remains largely unchanged, displays only a slightly larger splitting, suggesting that the magnetic field generated by the sample, $B_{z}^{\mathrm{s}}\approx-H$. Indeed, the extracted local field at the NV, $B_{z}=2.5~{}$G $\ll H=15~{}$G, corroborating the presence of superconducting Meissner phase that expels the external field. We emphasize that the observation of a near-complete cancellation of $H$ underscores the close proximity of our NV sensors to the superconducting BSCCO.
@@
As we increase the temperature, the near-complete expulsion of the external field persists $T\sim 40~{}$K. Beyond this point, ODMR spectrum splits into a characteristic four resonances, indicating that the field is no longer fully canceled and that the system has transitioned into a vortex phase, allowing magnetic flux to penetrate the sample. As the temperature further rises, the vortex density increases, raising the local field, $B_{z}$ until it reaches the applied field $H$ at $T=T_{c}\approx 90~{}$K, Interestingly, right below $T_{c}$, the local field $B_{z}$ is measured to be slightly greater than $H$, which may be attributed to a paramagnetic Meissner effect geim1998paramagnetic .
@@
Leveraging the local nature of our sensor, we then investigate the spatial dependence of the Meissner suppression. Performing ODMR spectroscopy along a line-cut across the BSCCO [black dashed line in Fig. 2(b)], we directly extract the local magnetic field $B_{z}$ as a function of spatial position at $T=30$ K [Fig. 2(c)]. Beneath the sample, the Meissner suppression is clearly observed at small applied field $H\lesssim 20~{}$G. However, when $H$ exceeds 25 G, which is on the same order as the independently characterized lower critical field $H_{\mathrm{c1}}$ (see Methods), the Meissner effect is suppressed, and $H$ is only partially expelled. Crucially, the residual field also exhibits spatial variations across the sample, indicating that magnetic flux penetrates the BSCCO, forming superconducting vortices. One such vortex, possibly pinned by a local defect, is imaged in Figure 2(d), where the locally measured magnetic field $B_{z}$ exceeds the applied field $H=40~{}$G due to the bunching of magnetic field lines.
@@
Zero-field superconducting fluctuations — Employing the full capabilities of the NV center, we go beyond conventional static magnetic field measurements by performing $T_{1}$ relaxometry of the NV to study superconducting fluctuations [Fig. 1(a)]. To this end, we first optically polarize the NV spin state to $|m_{s}=0\rangle$ and subsequently measure the timescale $T_{1}$ over which the spin relaxes back to thermal equilibrium [Fig. 3(a)]. Although the intrinsic relaxation of NV centers is primarily driven by spin-phonon interactions within the diamond itself, additional magnetic fluctuations from BSCCO, resonant with the NV’s bare frequency ($D_{\mathrm{gs}}=2.87~{}$GHz), can expedite the spin relaxation process and reduce the $T_{1}$ of NV.
@@
This is indeed borne out by our data. As shown in Figures 3(a)(b), NV centers located beneath the BSCCO exhibit shorter $T_{1}$ compared to NV centers situated away from the BSCCO across the entire temperature range covering the superconducting phase, criticality and metallic phase. This trend is further highlighted in Figure 3(c), which shows that the NV relaxation rate ($1/T_{1}$) as a function of a spatial line-cut across three distinct temperature regimes, with faster relaxation rates observed in all cases for NV centers beneath the BSCCO sample.

### v2

# Quantum noise spectroscopy of superconducting dynamics in thin film Bi2Sr2CaCu2O8+δ

@@ Abstract
Characterizing the low-energy dynamics of quantum materials is crucial to our understanding of strongly correlated electronic states. Yet, it remains experimentally challenging to investigate such dynamics with high spectroscopic resolution in both frequency and momentum space, particularly in two-dimensional correlated systems. Here, we leverage Nitrogen-Vacancy (NV) centers in diamond as a powerful and non-invasive tool to study thin-film Bi2Sr2CaCu2O8+δ (BSCCO), revealing several distinct dynamical phenomena across the superconducting phase diagram. At zero magnetic field and low temperatures, NV depolarization ($T_{1}$) noise spectroscopy captures the low-frequency (GHz-scale) magnetic noise generated by nodal superconducting quasiparticle excitations, in agreement with Bardeen–Cooper–Schrieffer (BCS) mean-field theory. Near the critical temperature $T_{c}\approx 90$ K, supercurrent-fluctuation-induced noise leads to a sharp reduction of the NV $T_{1}$. By carefully analyzing the temperature-scaling of $T_{1}$, we observe clear deviations from the BCS prediction, reflecting the importance of order parameter fluctuations and enabling the determination of both static and dynamical critical exponents. When a small field is applied, we detect a broad and asymmetric reduction of NV $T_{1}$ near $T_{c}$; the field-induced smearing of the transition unveils the presence of a vortex liquid phase. Finally, NV decoherence ($T_{2}$) noise spectroscopy allows us to characterize magnetic noise at even lower MHz-scale frequencies and obtain evidence for complex vortex-solid fluctuations well below $T_{c}$. Our results establish quantum noise spectroscopy as a versatile platform for probing dynamical phenomena in superconductors, with frequency and length scales complementary to existing techniques.
@@
Introduction — Strong electronic correlations are responsible for the plethora of complex emergent phenomena observed in quantum materials. This is perhaps best exemplified in unconventional superconductors, where electronic interactions induce a macroscopic quantum coherence of electron-pairs and dissipationless flow of current [70, 37, 75]. Key features of this state are encoded in the low-energy dynamics, manifesting as quasiparticle excitations arising from the breaking of electron-pairs [4, 5, 10], slow collective modes such as critical pairing fluctuations near the metal–superconductor phase transition [46, 65, 53, 14, 66], or topological excitations such as vortices carrying quantized magnetic flux [70, 9, 29, 7, 58]. Characterizing these dynamics thus offers a new avenue towards understanding the superconducting state. This is particularly important when addressing the emerging landscape of correlated, two-dimensional superconducting materials [3, 42], where the application of conventional experimental probes is a notorious challenge.
@@
Figure 1: NV noise spectroscopy of superconducting dynamics. (a) Summary of different experimental techniques for probing superconductivity, organized by their frequency and lengthscales, highlighting the complementarity of our qubit-based noise spectroscopy (NV $T_{1}$ and $T_{2}$) and DC magnetometry [6]. Techniques depicted include: electrical transport [22], superconducting quantum interference device (SQUID) [40, 20], Lorentz microscopy [30, 67], X-ray/neutron scattering [12, 57], magnetic resonance with nuclear, electronic, or muonic spins (NMR/ESR/µSR) [34, 31], angle-resolved photoemission spectroscopy (APRES) [51, 41], low-energy and transmission electron microscopy (LEEM/TEM) [2, 56], and scanning tunneling microscope (STM) [60]. Some techniques, while capable of accessing high momentum data, require macroscopic sample volumes and, thus, are only suitable for probing mm-scale samples. Techniques like pump-probe spectroscopy [36, 16], which are known to drive the sample strongly out of equilibrium, are not included. (b) Schematic of the experimental set-up: a thin film ($\sim 200$ nm thickness) of BSCCO is directly exfoliated onto a diamond plate containing an ensemble of NV centers within a thin layer ($\sim 50$ nm) beneath its surface. Current fluctuations in the BSCCO sample induce magnetic field noise at the NV locations which can be detected by either depolarization ($T_{1}$) or decoherence ($T_{2}$) noise spectroscopy. An out of plane magnetic field $H$ is applied, which is parallel to the quantization axis of the [111] NV group. The NV exhibits a splitting of its $|m_{s}=\pm 1\rangle$ sublevels, which is determined by the local magnetic field $B_{z}$, centered around a zero-field splitting $D_{\mathrm{gs}}=2.87~{}$GHz. (c) Schematic H-T phase diagram of a typical type-II superconductor [70, 37, 75]. The green arrows indicate zero-field cooling ($H=0$) and field-cooling ($H>0$) pathways used in our experiments.
@@
Over the past decades, several experimental techniques have been developed to probe correlated quantum dynamics in superconductors [12, 36, 16, 60, 22, 34]. Broadly, such techniques may be categorized into scattering methods, local spectroscopic probes, and bulk measurements [Fig. 1(a)]. Scattering probes (e.g. X-ray scattering [12]) provide both energy- and momentum-resolved data, but they typically require experimental apparatus beyond table-top settings, large sample volume, and can drive the system out of equilibrium [36, 16]. Local probes such as scanning tunneling microscopy (STM) [60] are applicable to low-dimensional samples and provide lattice-scale spatial resolution, yet they exhibit limited momentum resolution and typically operate at low temperatures. By contrast, bulk measurements (including transport [22] and nuclear magnetic resonance [34]) can operate over a wide temperature range, but generally demand large sample volume. Therefore, resolving spectral features at long wave-lengths below the meV scale in two-dimensional materials and thin film samples remains a challenge, leading to an unfulfilled gap for experimental techniques that can probe equilibrium dynamics at both low-energies and a variety of length-scales.
@@
In this work, we fill this gap by introducing a non-invasive, table-top experimental technique — quantum noise spectroscopy — to probe a wide range of low-energy dynamical phenomena in superconductors. Specifically, by using an ensemble of nitrogen-vacancy (NV) centers positioned in close proximity to a thin film high-$T_{c}$ cuprate — Bi2Sr2CaCu2O8+δ (BSCCO, Bi-2212), we perform noise spectroscopy of superconducting fluctuations as a function of temperature, frequency and applied magnetic field. We demonstrate in-situ detection of spatially resolved dynamics spanning a broad frequency range (MHz to GHz) and length-scales (nm to $\mathrm{\mu m}$), thus offering a powerful complementary toolset to existing techniques for probing superconductivity [Fig. 1(a)].
@@
Our results are threefold. First, at zero applied field, we observe and characterize the fluctuations arising from both long-lived quasiparticle excitations at low temperature, and collective critical modes near the metal-superconductor transition temperature $T_{c}$. These fluctuations induce GHz-scale magnetic noise at the location of the nearby NVs which can be characterized via NV depolarization ($T_{1}$) spectroscopy. These capabilities enable us to directly measure the characteristic timescale $\tau$ of critical slowing down of pairing fluctuations near $T_{c}$ and determine both the static and dynamical critical exponents associated with the phase transition [45, 44, 33].
@@
Second, we directly measure the broadening of the transition in the presence of a small applied field $H$. At non-zero $H$, the system also hosts superconducting vortices whose dynamics induce additional magnetic field fluctuations. The field dependence of the NV relaxation rate is consistent with the existence of an intermediate-temperature vortex liquid phase [9, 29], where magnetic flux lines diffuse through the superconducting film.
@@
Finally, we employ decoherence ($T_{2}$) noise spectroscopy to study lower-frequency fluctuations in the MHz regime [52]. Deep within the superconducting phase, we observe a distinct noise mechanism that significantly enhances the NV decoherence rate. These strong, low-frequency magnetic fluctuations potentially arise from thermally assisted flux motion in a vortex solid phase [9, 24, 23].
@@
Figure 2: Observing Meissner effect and superconducting vortices. (a) ODMR spectrum of NV centers under BSCCO with an applied field $H=15$ G at temperatures from 10 K to 120 K. The colors represent the normalized fluorescence contrast (FL) of the spectrum. The gray dashed line marks the superconducting transition temperature $T_{c}\approx 90$ K. Top: at 10 K, ODMR spectrum of NV centers under BSCCO at $H=0$ (yellow), where the intrinsic splitting $2\Pi_{\perp}$ originates from the local electric field [55]. At $H=15$ G, the spectrum (blue) remains largely unchanged, indicating a strong magnetic field repulsion due to the Meissner effect. Bottom: when the bias field is not suppressed (120 K), the [111] NV group aligned with the bias field exhibits a splitting $\Delta$, while the remaining NVs in other three crystal orientations are degenerate, contributing to the inner two peaks. (b) Optical image of the thin film BSCCO exfoliated on a diamond surface. (c) Local magnetic field $B_{z}$ as a function of the applied field $H$ along a line-cut across the BSCCO [black dashed line in (b)] at 30 K. Compared to NV centers far away from BSCCO, those under BSCCO (gray region) experience smaller local magnetic fields due to the Meissner suppression. Near the edge of the BSCCO, $B_{z}$ gets enhanced due to the concentration of the magnetic field lines. (d) Schematic of magnetic flux penetrating a superconductor in its vortex phase. Top: local field image of a defect-trapped superconducting vortex at $H=40$ G and $T=40$ K. (e) Schematic of magnetic flux being repelled by a superconductor in its Meissner phase.
@@
Experimental setup — For our experiments, we study Bi2Sr2CaCu2O8+δ (BSCCO), a prominent type-II superconductor with a layered crystal structure and a high critical temperature $T_{c}$ exceeding the boiling point of liquid nitrogen [37, 75, 9]. This compound exhibits a rich and complex phase diagram, making it an ideal platform for investigating a variety of different types of low-energy dynamics [Fig. 1(c)]. Bulk crystals of nearly optimally doped BSCCO are grown by the traveling floating zone method [71], with $T_{c}\approx 90~{}$K measured via transport (see Methods). To perform noise spectroscopy with NV centers, we exfoliate the BSCCO to a 200200200~{}200nm thick film and transfer it directly onto a diamond sample containing an ensemble of NV centers in a thin layer $\sim 50$ nm beneath the diamond surface [Fig. 1(b)].
@@
Each NV center hosts a spin-1 electronic ground state, with Hamiltonian [63]:
@@
where {$S_{x}$, $S_{y}$, $S_{z}$} are the spin-1 operators of the NV with the N–V axis defining the quantization axis ($\hat{z}$), $D_{\mathrm{gs}}=2.87~{}$GHz is the zero-field splitting between $|m_{s}=0\rangle$ and $|m_{s}=\pm 1\rangle$ sublevels, $\Pi_{\perp}=\sqrt{\Pi_{x}^{2}+\Pi_{y}^{2}}=6.6~{}$MHz characterizes the coupling to local electric fields [55], $B_{z}$ is the local magnetic field along the NV axis, and $\gamma_{e}$ is the electronic spin gyromagnetic ratio.
@@
The NV energy levels can be probed via optically detected magnetic resonance (ODMR) spectroscopy where one detects a decrease in NV fluorescence intensity when sweeping the frequency of an applied microwave through the resonances $|m_{s}=0\rangle\leftrightarrow|m_{s}=\pm 1\rangle$. In the experiment, we use a [111]-cut diamond crystal and apply a tunable magnetic field $H$ along the out-of-plane direction [Fig. 1(b)]. In this geometry, the NV crystallographic group aligned with the applied field exhibits the largest splitting, $\Delta=\sqrt{(2\Pi_{\perp})^{2}+(2\gamma_{e}B_{z})^{2}}$ between $|m_{s}=\pm 1\rangle$ sublevels, where $B_{z}$ includes both contributions from the applied field $H$ and the $B^{\mathrm{s}}_{z}$ generated from the superconducting sample ($B_{z}=H+B_{z}^{\mathrm{s}}$). The other three NV groups are always observed to be degenerate.
@@
Imaging the Meissner effect and flux vortices — Using the NV centers as sensitive local magnetometers, we begin by determining the presence of superconductivity in BSCCO via the DC Meissner effect [64, 74, 48, 8, 59, 32]. After zero-field cooling the sample down to 101010~{}10K (well below $T_{c}$), we perform an ODMR measurement with $H=0$ [Fig. 2(a)]. The resulting spectrum shows two resonances separated by the intrinsic splitting $2\Pi_{\perp}$. When we apply a small field $H=15~{}$G, the ODMR spectrum remains largely unchanged, displaying only a slightly larger splitting which suggests that the magnetic field generated by the sample cancels the applied one, $B_{z}^{\mathrm{s}}\approx-H$. Indeed, the extracted local field at the NV, $B_{z}=2.5~{}$G $\ll H=15~{}$G, corroborating the presence of superconducting Meissner phase that expels the external field. We emphasize that the observation of a near-complete cancellation of $H$ implies a close proximity of our NV sensors to the superconducting sample.
@@
As we increase the temperature, the near-complete expulsion of the external field persists until $T\sim 40~{}$K. Beyond this point, ODMR spectrum splits into a characteristic four resonances, indicating that the field is no longer fully canceled and that the system has transitioned into a vortex phase, allowing magnetic flux to penetrate the sample. As the temperature further rises, the vortex density increases, raising the local field, $B_{z}$ until it reaches the applied field $H$ at $T=T_{c}\approx 90~{}$K. Interestingly, right below $T_{c}$, the local field $B_{z}$ is measured to be slightly greater than $H$, which may be attributed to a paramagnetic Meissner effect [25].
@@
Leveraging the local nature of our sensor, we then investigate the spatial dependence of the Meissner suppression. Performing ODMR spectroscopy along a line-cut across the BSCCO sample [black dashed line in Fig. 2(b)], we directly extract the local magnetic field $B_{z}$ as a function of spatial position at $T=30$ K [Fig. 2(c)]. Beneath the sample, the Meissner suppression is clearly observed at small applied field $H\lesssim 20~{}$G. However, when $H$ exceeds 25 G, which is on the same order as the independently characterized lower critical field $H_{\mathrm{c1}}$ (see Methods), the Meissner effect is suppressed, and $H$ is only partially expelled. Crucially, the residual field also exhibits spatial variations across the sample, indicating that magnetic flux penetrates the BSCCO, forming superconducting vortices. One such vortex, possibly pinned by a local defect, is imaged in Fig. 2(d), where the locally measured magnetic field $B_{z}$ exceeds the applied field $H=40~{}$G due to the bunching of magnetic field lines.
Figure 3: Zero-field superconducting fluctuations. (a) $T_{1}$ decay profile for NV centers located under BSCCO (left, blue) and away from BSCCO (right, orange) at 110, 93 and 40 K. Inset: Location of the measured NV centers with respect to the BSCCO sample. (b) The $T_{1}$ relaxation timescales as a function of temperature measured on the NV centers located under and away from BSCCO. (c) Spin relaxation rate, $1/T_{1}$, along a spatial line-cut across the BSCCO sample (see inset) at 110, 93 and 40 K. The relaxation rate increases for NV centers under BSCCO (shadow region), particularly near the critical temperature $T_{c}$. (d) The BSCCO-induced decay rate of NVs, $\Gamma_{1}(T)$, as a function of temperature with $H=0$. Inset: a power-law relationship at low temperatures: $\Gamma_{1}(T)\propto T^{2}$, is consistent with a mean-field approach accounting for nodal quasiparticles in d-wave superconductors. (e) $\Gamma_{1}(T)$ as a function of the absolute temperature deviation from the critical temperature, $|T-T_{c}|$, for both the metallic and superconducting (SC) sides of the transition. BCS mean-field quasiparticle dynamics (dashed line) predicts $\Gamma_{1}(T)\propto|T-T_{c}|^{-x}$ with $x=1/2$. However, our experimental data is consistent with $x\approx 1$, suggesting that additional superconducting order parameter fluctuations are important. This data aligns closely with a theoretical model of critical Langevin dynamics of the order parameter (solid lines).
@@
Zero-field superconducting fluctuations — Employing the full capabilities of the NV center, we go beyond conventional static magnetic field measurements and study superconducting fluctuations by performing $T_{1}$ relaxometry of the NV [Fig. 1(b)]. To this end, we first optically polarize the NV spin state to $|m_{s}=0\rangle$ and subsequently measure the timescale $T_{1}$ over which the spin relaxes back to thermal equilibrium [Fig. 3(a)]. Although spin-phonon interactions within the diamond itself lead to an intrinsic relaxation rate of NV centers, additional magnetic fluctuations from BSCCO, resonant with the NV’s bare frequency ($D_{\mathrm{gs}}=2.87~{}$GHz), can expedite the spin relaxation process and reduce the $T_{1}$ of NV.
@@
This is indeed borne out by our data. As shown in Figures 3(a)(b), NV centers located beneath the BSCCO exhibit shorter $T_{1}$ compared to NV centers situated away from the BSCCO across the entire temperature range. This trend is further highlighted in Figure 3(c), which shows that the NV relaxation rate ($1/T_{1}$) as a function of a spatial line-cut across three distinct temperature regimes — the superconducting phase, the critical region and the metallic phase, with faster relaxation rates observed in all cases for NV centers beneath the BSCCO sample.